Solve the following system of linear equation.
y=2x+1
x+y=-2

Answer by checkley71(8403)

y=2x-1 then x+2x-1=-2 or 3x=-2+1 or 3x=-1 or x=-1/3 then y=2(-1/3)-1 or
y=-2/3-1 or y=-1 2/3